Solving this following ODE by Frobenius Method
X^2y’’+(4x-x^2)y’+(2-x)y=0
1
Expert's answer
2019-05-13T03:17:09-0400
"x^2y''+(4x-x^2)y'+(2-x)y=0 \ \ \ \ \ \ \ \ \ (1)"

Dividing by "x^2"


"y''+{4x-x^2 \over x^2}y'+{2-x\over x^2}y=0""xP(x)=4-x, x^2Q(x)=2-x"

Both "P(x)" and "Q(x)" are analytic at "x=0." A singular point "x=0" is a regular singular point of the equation (1). Then we can find at least one solution to a second order differential equation by assuming a solution of the form: 


"y=\displaystyle\sum_{n=0}^\infin a_nx^{n+r}, a_0\not =0""y'=\displaystyle\sum_{n=0}^\infin a_n(n+r)x^{n+r-1}""y''=\displaystyle\sum_{n=0}^\infin a_n(n+r)(n+r-1)x^{n+r-2}"

Substitute


"x^2 \displaystyle\sum_{n=0}^\infin a_n(n+r)(n+r-1)x^{n+r-2}+(4x-x^2)\displaystyle\sum_{n=0}^\infin a_n(n+r)x^{n+r-1}+""+(2-x)\displaystyle\sum_{n=0}^\infin a_nx^{n+r}=0"

Or


"\displaystyle\sum_{n=0}^\infin a_n(n+r)(n+r-1)x^{n+r}+\displaystyle\sum_{n=0}^\infin a_n[4(n+r)+2]x^{n+r}+""+\displaystyle\sum_{n=1}^\infin a_{n-1}(-n-r)x^{n+r}=0"




"a_0r(r-1)x^r+a_0(4r+2)x^r=0""(r+2)(r+1)=0""r_1=-2, r_2=-1"




"a_1(1+r)rx^{r+1}+a_1(4(1+r)+2)x^{r+1}+a_0(-1-r)x^{1+r}=0""a_1(r+2)(r+3)x^{r+1}=a_0(1+r)x^{1+r}"

"a_2(2+r)(1+r)x^{r+2}+a_2(4(2+r)+2)x^{r+2}+a_1(-2-r)x^{r+2}=0""a_2(r+3)(r+4)x^{r+2}=a_1(2+r)x^{2+r}"


"a_3(3+r)(2+r)x^{r+3}+a_3(4(3+r)+2)x^{r+3}+a_2(-3-r)x^{r+3}=0""a_3(r+4)(r+5)x^{r+3}=a_2(3+r)x^{3+r}"


"a_n(r+n+1)(r+n+2)x^{r+n}=a_{n-1}(n+r)x^{n+r}"

"r_1=-2""a_1(-2+2)(-2+3)x^{-2+1}=a_0(1-2)x^{1-2}, a_0=0"


"r_2=-1"

"a_1(-1+2)(-1+3)x^{-1+1}=a_0(1-1)x^{1-1}, a_1=0, a_0\not =0"

"a_2(-1+3)(-1+4)x^{-1+2}=a_1(-1+r)x^{-1+r}, a_2=0"


"a_2=a_3=...=a_n=...=0"

"\widetilde{y_2}=a_0x^{0-1}={a_0 \over x}"

"y={a_0 \over x}, a_0\in \R, a_0\not= 0"